An archer shoots an arrow at a 75.0 m distant target, the bull's-eye of which is at same height as the release height of the arrow. (a) At what angle must the arrow be released to hit the bull's-eye if its initial speed is 36.0 m/s? (Although neglected here, the atmosphere provides significant lift to real arrows.) ° (b) There is a large tree halfway between the archer and the target with an overhanging branch 3.50 m above the release height of the arrow. Will the arrow go over or under the branch? under over
Answer #1

a)
I think that you can use the range equation to figure out the angle (θ):
R=\frac{[(V_o)^2 \times 2 sin \theta ]}{g}
Where

R=75.0m \\ Vo=36m/s \\ g=9.8m/s^2

So for θ:

75= \frac{{(36)}^2 \times 2 sin \theta }{9.8}
75 \times 9.8=[{(36)}^2 \times 2Sin \theta ] \\ 735=1296 \times 2Sin \theta \\ \frac{735}{2 \times 1296}=Sin \theta\\ 0.284=Sin \theta\\ Sin^{-1}{(0.284)}= \theta\\ \theta =16.49º

b.)
Find time (t), the distance that the arrow takes to travel \frac{75}{2}=37.5m
Vx=Vo \times Cos16.02º \\ Vx=36m/s \times Cos16.02º \\ Vx=34.6m/s
D=Vx \times t


Solving for t:

37.5=9.94 \times t\\ t=\frac{37.5}{34.6}\\ t=1.08s

Vy=VoSin16.02º \\ Vy=36.0 Sin16.02º \\ Vy=9.94m/s
Use the following equation and solve for d:
d=V_y \times t+1/2(-g)t^2
Where

V_y=9.94m/s \\ g=9.8m/s \\ t=1.08s

So,

d=(9.94 \times 1.08)+1/2(-9.8)(1.08)^2\\ d=10.74m-5.7m \\ d=5.03m

Answer.